OpenStudy (anonymous):

find the exact value of the trigonometric function at the given real number.If the value is undefined, enter no solution. sin17(pie) cos19(pie)/2 tan13(pie)/2

OpenStudy (anonymous):

good god

OpenStudy (anonymous):

first of all sine is defined for all real numbes, so it cannot be undefined

OpenStudy (anonymous):

second of all the symbol \[\pi\] is a greek letter, if you want to write it in english it is called "pi" not "pie"

OpenStudy (cwrw238):

sin 17pi = sin pi = 0 i think!!!

OpenStudy (anonymous):

if you want to find the value of \[\sin(17\pi)\] you can say that since sine is periodic with period \[2\pi\] then \[\sin(17\pi)=\sin (\pi)=0\]

OpenStudy (cwrw238):

ah